# Contrastive-KAN: A Semi-Supervised Intrusion Detection Framework for Cybersecurity with scarce Labeled Data

### v1

@@ 4.4 Fine-tuning
After pre-training the $KAN_{\theta}(\cdot)$, we stack another KAN layer, denoted as $KAN_{\lambda}(\cdot)$, with learnable parameters $\lambda$, to form the final classifier. From this point onward, we no longer use batches in the notation. One important aspect of the training process is that only the new layer parameters, denoted by $\lambda$, are updated, while the pre-trained parameters $\theta$ remain frozen, and will not be updated in the fine-tuning stage. The labeled dataset used during the fine-tuning is denoted as $D=(X_{scaled},Y)$, where $X_{scaled}$ is the standardized input feature vector and $Y^{i}\in\{0,\dots,C\}$ is the label. Unlike the pre-training phase, that the labels were unknown, in fine-tuning stage the labels are used during training. The notation for the $i^{th}$ data point is defined as:
$z^{i}=KAN_{\theta}(X_{scaled}^{i})$ (17)
@@
$class^{(i)}=softmax(\zeta^{(i)})$ (19)
$error_{final}=\sum_{i=1}^{N}crossEntropy(class^{(i)},Y^{(i)})$ (20)
First we obtain the latent space representations from the feature extractor $KAN_{\theta}(\cdot)$, as defined in Equation 17. These representations are then passed through the newly stacked KAN layer $KAN_{\lambda}(\cdot)$ to produce the class representations, based on Equation 18. These class representations are then passed to the softmax function, which will form the class probabilities. Next, the class probabilities are fed into the cross-entropy loss function, in Equation 20 which will create the final classification loss, $error_{final}$. This loss will be used in gradient descent algorithm to update the model parameters according to:
$\lambda_{n+1}=\lambda_{n}-\alpha\frac{\partial error_{final}}{\partial\lambda}$ (21)
where $\alpha$ is the fine-tuning learning rate, and $\lambda$ is the final layer’s parameters. After completing the training process, the model can be saved and used for inference on new data. Algorithm 2, shows a pseudo-code of the fine-tuning phase of the proposed method.
@@
1:  Normalize with the $\mu$ and $\sigma$ derived from pre-training:$X_{\text{scaled}}\leftarrow\frac{X-\mu}{\sigma}$
@@
6:     Feed the data point $X^{i}_{scaled}$ $KAN_{\theta}(\cdot)$:$z^{i}\leftarrow KAN_{\theta}(X_{\text{scaled}}^{i})$
7:     Get class representations:$\zeta^{(i)}\leftarrow KAN_{\lambda}(z^{i})$
8:     Class probabilities: $class^{(i)}\leftarrow softmax(\zeta^{(i)})$
9:     Compute classification error: $error_{final}\leftarrow\sum_{i=1}^{N}crossEntropy(class^{(i)},Y^{(i)})$
10:     Update final KAN layer $KAN_{\lambda}(\cdot)$ parameters:$\lambda_{n}\leftarrow\lambda_{n-1}-\alpha\frac{\partial error_{final}}{\partial\lambda}$
@@ 4.5 Simulation Setup
The implementation of this project is based on efficient-kan implementation111www.github.com/Blealtan/efficient-kan. In this experiments, we utilize a learning rate of 0.001 for both pre-training and fine tuning stages across all datasets. The contrastive pre-training batch sizes are 64, 16, and 64 for the UNSW-NB15, BoT-IoT and Gas Pipeline datasets, respectively. And likewise, for fine-tuning, the batch sizes are 64, 32, and 2, respectively. The masking percentage is set to $10\%$ for UNSW-NB15 dataset and $20\%$ for BoT-IoT and Gas Pipeline dataset. The masking percentage for the BoT-IoT and Gas Pipeline dataset is higher, due to their lower feature dimensionality. With a smaller $p$, there is a chance of similar augmented views after performing the masking. The Adam optimizer is used during the pre-training phase, while AdamW is used for fine-tuning stage. The experiments are performed on Kaggle environment, with 32 GB of RAM, an NVIDIA Tesla P100 GPU with 16 GB of VRAM, and a Intel(R) Xeon(R) CPU 2.00 GHz.
@@ 5.1 Performance Analysis
In this section, we analyze the results of the experiments on the three different datasets, described earlier. Figure 3 presents the t-SNE visualizations of the test data passed through the KAN feature extractor before pre-training and Figure 4 shows the corresponding results after completing the pre-training. These two figures demonstrate the effectiveness of the proposed contrastive learning framework. In the BoT-IoT dataset, the DoS and DDoS samples form more compact clusters and exhibit reduced overlap after the network has been pre-trained. In the UNSW-NB15 dataset, after the pre-training, the Exploits and Normal samples demonstrate lower degree of overlap, and we witness better separability between Normal and attack samples. Additionally, the Exploits and Fuzzers data points are better clustered in comparison to non-pretrained FE. The t-SNE visualizations for gas pipeline clearly confirm the positive effect of the contrastive pre-training on the separability of the test data. However, the gas pipeline dataset is the most challenging of all three, due to complex structure, smaller volume of data and fewer distinguishable features.
Figure 5: Visualization of the spline-based learned activation functions, aggregated with SiLU, for UNSW-NB15 dataset. The red points represent the control points that were mentioned in Equation 6. The observed upward trend in the curves results from the addition of SiLU function to the spline components, as seen in Equation 4.
@@ 5.2 Inference Time
We measured the inference time of the network to assess its feasibility for real-time deployment across all of datasets. We evaluated the proposed contrastive-KAN model on the entire test set and reported the average inference time per sample. For Gas Pipeline dataset the inference time was 0.41 ms; for the UNSW-NB15 dataset, it was 0.26 ms; and for BoT-IoT it was 0.13 ms. All evaluations are performed on CPU from the Kaggle environment, with the batch sizes specified during fine-tuning stage. It is notable that smaller batch sizes introduce some computational overhead, because smaller batch sizes require a larger number of iterations to process the entire dataset. With this said, the evaluation times clearly demonstrate that our proposed method is real-time, providing fast and timely intrusion detection.
@@ 5.3 Model Interpretability
Since the need for safe and reliable operation in the IIoT and IoT networks is present, explainable artificial intelligence (XAI) and model interpretability have become essential. For this purpose, KAN provides a very valuable aspect, that allows the visualization of the leaned activation functions. With this feature, engineers can extract rules form the trained intrusion detection systems, which can then be used in the rule-based approaches. This will further insure the reliability and safety of operation of the systems especially in safety-critical applications.
In Figure 5, the spline functions learned by the pre-trained FE of the UNSW-NB15 dataset, biased by SiLU activation function, are visualized. The red dots, represent the control points, discussed in Equation 6. The upward trend is due to the addition of the SiLU activation function. While there are many leaned activation functions, it’s clear that visualizing all of them is not possible. Figure 5 demonstrates the feasibility of the mentioned feature.
Table 4: UNSW-NB15 f1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with dimensions [204 50 6].
@@
Table 5: BoT-IoT f1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with dimensions [20 100 2].
@@
Table 6: Gas Pipeline F1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with input [19 100 8].
@@
Table 7: Ablation study on the UNSW-NB15 dataset using the specified data split. The grid size and spline order are derived from the best results of Table 4.
@@
Table 8: Ablation study on the BoT-IoT dataset using the specified data split. The grid size and spline order are derived from the best results of Table 5.
@@
Table 9: Ablation study on the Gas Pipeline dataset, using the specified data split. The grid size and spline order are derived from the best results of Table 6.
@@ 5.4 Ablation Study
For the ablation study, we performed experiments focusing on two key KAN hyper-parameters: spline order and grid size. While KAN has multiple hyper-parameters, spline order and grid size are the most influential. The results in Table 4 and Table 5 indicate that a network with higher grid size and smaller spline order manages to outperforms the cases with higher spline order. This is exactly what is mentioned in [52], which states that layers with higher grid sizes perform more robustly, despite having a higher Lipchitz constant. However, this trend does not hold for the Gas Pipeline dataset in Table 6.
Similarly we examined the performance of KAN against MLP, in two training paradigms: contrastive pre-training and end-to-end training. In Table 7-Table 9, the experiments conducted are listed. For each dataset we perform 12 experiments to determine the effect of the training paradigm and the width of the network. In these tables, we use two-layer KANs and three-layer MLP networks. For KANs, the spline order and grid size parameters are chosen from the best result of Table 4-Table 6. Under architecture columns, the first and last numbers represent the dimensionality of the input data and the number of classes that the classification is being performed on, and in MLP cases, the middle numbers indicate the dimensionality of the middle layer. And beside the architecture, the number of parameters of each network is present. For MLP, the focus is on an architecture that maintains a parameter count as in its KAN counterpart. All models, whether trained contrastively or end-to-end, are trained using the specified data splits. The difference is that, in the end-to-end setting, only the fine-tuning set is used. It is observed that KAN’s performance does not necessarily improve proportionate to the parameter count. In contrary, for MLP networks, the performance improves with increase in the number of parameters. One can deduce that tuning the hyper-parameters of KAN is much difficult compared to the parameters of MLP networks. Although the Gas Pipeline dataset is very challenging, the parameter count of models trained on Gas Pipeline is significantly lower than UNSW-NB15 and BoT-IoT.
@@
It is observable that the KAN outperforms the MLP networks, that are tailored to have the same number of parameters as the KAN. In the end-to-end training, the networks are trained using the same fine-tuning set, and the pre-training with unlabeled data is not performed. The contrastive training, taking advantage of the large volumes of unlabeled data, further improves the performance.
@@
It is worth noting that, for KANs, the width of the network has little impact on the performance. As can be observed, the networks with width of 50 or 100, outperforms cases with widths 150 and 200. This supports the earlier findings that the main factors influencing the performance of KAN are the grid size and spline order. However, for MLP networks, increasing the width of the networks, often results in marginally better performance. This indicates that, for MLP networks, there is a more predictable relationship between network capacity and the performance, or in other words, MLP networks are easier to tune.
@@
Table 10: Comparison of our proposed methods with other approaches, UNSW-NB15 dataset.
@@
Table 11: Comparison of our proposed methods with other approaches, BoT-IoT dataset.
@@ 5.5 Comparative Analysis
Table 10, Table 11 and Table 12 show a comparative analysis of the state-of-the-art. The performance of the propose method is superior compared to other learning-based algorithms, some of which utilize more labeled data than the proposed method. This proves that the proposed algorithm is a good suit for scenarios with label scarcity.
@@ 6 Conclusion
In this paper, we proposed a semi-supervised contrastive learning approach for real-time intrusion detection under limited labeled data conditions. This framework demonstrate strong performance on IoT and IIoT datasets, outperforming other state-of-the-art methods. The results produced by this study shows that the performance of the KAN is superior to that of the MLP networks, when labeled data is scarce. Additionally, the contrastive training improves the performance of the network by utilizing the available unlabeled data. A key advantage of the proposed method is the the interpretability provided by KAN, in form of spline function visualization. This may be utilized to extract relevant rules from the network, which is beneficial in safety-critical applications, that further increases reliability, stability, and ensuring safe operation. One of the findings of this paper is that the spline order and grid size are the most influential KAN hyper-parameters to tune, while the layer dimensions have minimal impact on improving the performance. For our datasets, a higher grid size and lower spline order yields better performance. The other key finding, is that the hyper-parameter tuning for MLP networks are much easier compared to their KAN counterpart, since KAN does not exhibit a straightforward relationship between performance and the number of parameters or the increase in spline order and grid size. Although this study performed well in intrusion detection problems with limited labeled data, further studies must be performed on KAN architecture for an unsupervised case, where the training requires no labeled data, which is left for future work.

### v2

@@ 4.4 Fine-tuning
After pre-training the $KAN_{\theta}(\cdot)$, we stack another KAN layer, denoted as $KAN_{\lambda}(\cdot)$, with learnable parameters $\lambda$, to form the final classifier. From this point onward, batch notation is omitted for simplicity. An important aspect of the training process is that only the new layer parameters, denoted by $\lambda$, are updated, while the pre-trained parameters $\theta$ remain frozen and are not updated during the fine-tuning stage. The labeled dataset used during fine-tuning is denoted as $D=(X_{\text{scaled}},Y)$, where $X_{\text{scaled}}$ is the standardized input feature vector and $Y^{i}\in\{0,\dots,C\}$ is the label. Unlike in the pre-training phase, where labels were unknown, the fine-tuning stage uses the labels during training. The notation for the $i^{th}$ data point is defined as:
$z^{i}=KAN_{\theta}(X_{\text{scaled}}^{i})$ (17)
@@
$class^{(i)}=\mathrm{softmax}(\zeta^{(i)})$ (19)
$\text{error}_{\text{final}}=\sum_{i=1}^{N}\text{crossEntropy}(class^{(i)},Y^{(i)})$ (20)
First, we obtain the latent space representations from the feature extractor $KAN_{\theta}(\cdot)$, as defined in Equation 17. These representations are then passed through the newly stacked KAN layer $KAN_{\lambda}(\cdot)$ to produce the class representations, based on Equation 18. The resulting class representations are fed into the softmax function to generate class probabilities. Next, these probabilities are used in the cross-entropy loss function, shown in Equation 20, which computes the final classification loss, $\text{error}_{\text{final}}$. This loss is then employed in a gradient descent algorithm to update the model parameters as follows:
$\lambda_{n+1}=\lambda_{n}-\alpha\frac{\partial\text{error}_{\text{final}}}{\partial\lambda}$ (21)
where $\alpha$ is the fine-tuning learning rate, and $\lambda$ denotes the parameters of the final layer. After completing the training process, the model can be saved and used for inference on new data. Algorithm 2 presents the pseudo-code for the fine-tuning phase of the proposed method.
@@
[…]
@@
6:  Feed the data point $X^{i}_{scaled}$ to $KAN_{\theta}(\cdot)$:$z^{i}\leftarrow KAN_{\theta}(X_{\text{scaled}}^{i})$
[…]
8:  Class probabilities: $class^{(i)}\leftarrow\mathrm{softmax}(\zeta^{(i)})$
9:  Compute classification error: $\text{error}_{\text{final}}\leftarrow\sum_{i=1}^{N}\text{crossEntropy}(class^{(i)},Y^{(i)})$
[…]
@@ 4.5 Simulation Setup
[…]
@@ 5.1 Performance Analysis
In this section, we analyze the results of the experiments conducted on the three different datasets described earlier. Figure 3 presents the t-SNE visualizations of the test data passed through the KAN feature extractor before pre-training, while Figure 4 shows the corresponding results after completing the pre-training. These figures demonstrate the effectiveness of the proposed contrastive learning framework.
In the BoT-IoT dataset, the DoS and DDoS samples form more compact clusters and exhibit reduced overlap after the network has been pre-trained. In the UNSW-NB15 dataset, following pre-training, the Exploits and Normal samples demonstrate a lower degree of overlap, resulting in better separability between Normal and attack samples. Additionally, the Exploits and Fuzzers data points are more distinctly clustered compared to the non-pretrained feature extractor. The t-SNE visualizations for the Gas Pipeline dataset clearly confirm the positive effect of contrastive pre-training on the separability of the test data. However, the Gas Pipeline dataset remains the most challenging among the three, due to its complex structure, smaller data volume, and fewer distinguishable features.
Figure 5: Visualization of the spline-based learned activation functions, aggregated with SiLU, for the UNSW-NB15 dataset. The red points represent the control points referenced in Equation 6. The observed upward trend in the curves results from the addition of the SiLU function to the spline components, as described in Equation 4.
@@ 5.2 Inference Time
We measured the inference time of the network to assess its feasibility for real-time deployment across all datasets. The proposed contrastive-KAN model was evaluated on the entire test set, and the average inference time per sample was reported. Specifically, the inference times were 0.41 ms for the Gas Pipeline dataset, 0.26 ms for the UNSW-NB15 dataset, and 0.13 ms for the BoT-IoT dataset. All evaluations were conducted on a CPU in the Kaggle environment, using the batch sizes specified during the fine-tuning stage. It is worth noting that smaller batch sizes introduce some computational overhead, as they require more iterations to process the entire dataset. Despite this, the measured inference times clearly demonstrate that our proposed method operates in real-time, providing fast and timely intrusion detection.
@@ 5.3 Model Interpretability
Given the critical need for safe and reliable operation in IIoT and IoT networks, explainable artificial intelligence (XAI) and model interpretability have become essential. In this context, KAN offers a valuable advantage by enabling the visualization of the learned activation functions. This feature allows engineers to extract meaningful rules from the trained intrusion detection system, which can subsequently be incorporated into rule-based approaches. Such interpretability further enhances the reliability and safety of these systems, particularly in safety-critical applications.
In Figure 5, the spline functions learned by the pre-trained feature extractor on the UNSW-NB15 dataset, biased by the SiLU activation function, are visualized. The red dots represent the control points, as discussed in Equation 6. The observed upward trend in the curves results from the addition of the SiLU activation function. Although many activation functions are learned, visualizing all of them is impractical. Nonetheless, Figure 5 effectively demonstrates the feasibility of this interpretability feature.
Table 4: F1-scores on the UNSW-NB15 dataset for various combinations of spline orders and grid sizes, obtained by training with the proposed framework. The network architecture consists of two layers with dimensions [204, 50, 6].
@@
Table 5: F1-scores on the BoT-IoT dataset for various combinations of spline orders and grid sizes, obtained by training with the proposed framework. The network architecture consists of two layers with dimensions [20, 100, 2].
@@
Table 6: F1-scores on the Gas Pipeline dataset for various combinations of spline orders and grid sizes, trained using the proposed framework. The network consists of two layers with input dimensions [19, 100, 8].
@@
Table 7: Ablation study on the UNSW-NB15 dataset using the specified data split. The grid size and spline order are selected based on the best results reported in Table 4.
@@
Table 8: Ablation study on the BoT-IoT dataset using the specified data split. The grid size and spline order are selected based on the best results reported in Table 5.
@@
Table 9: Ablation study on the Gas Pipeline dataset using the specified data split. The grid size and spline order are chosen based on the best results in Table 6.
@@ 5.4 Ablation Study
For the ablation study, we focused on two key KAN hyperparameters: spline order and grid size. Although KAN includes several hyperparameters, spline order and grid size are the most influential. The results shown in Table 4 and Table 5 indicate that networks with higher grid sizes and smaller spline orders outperform those with higher spline orders. This observation aligns with the findings in [52], which suggest that layers with larger grid sizes tend to be more robust, despite exhibiting a higher Lipschitz constant. However, this trend does not hold for the Gas Pipeline dataset, as shown in Table 6.
Similarly, we examined the performance of KAN against MLP under two training paradigms: contrastive pre-training and end-to-end training. The experiments conducted are summarized in Table 7 through Table 9. For each dataset, we performed 12 experiments to evaluate the effects of the training paradigm and network width. In these tables, two-layer KANs and three-layer MLP networks are compared. For the KAN models, the spline order and grid size parameters were selected based on the best results from Table 4 to Table 6.
In the architecture columns, the first and last numbers denote the input dimensionality and the number of output classes, respectively. For MLPs, the middle number indicates the dimensionality of the hidden layer. Next to the architecture, the total number of parameters for each network is reported. For MLPs, the architectures were chosen to maintain a similar parameter count to their KAN counterparts.
All models, whether trained with contrastive pre-training or end-to-end, were trained using the specified data splits. The main difference is that in the end-to-end training paradigm, only the fine-tuning dataset was used.
It was observed that KAN’s performance does not necessarily improve proportionally with the number of parameters. In contrast, MLP networks tend to perform better as the number of parameters increases. This suggests that tuning KAN’s hyperparameters is more challenging compared to tuning the parameters of MLP networks. Despite the Gas Pipeline dataset being the most challenging, the parameter counts of models trained on it are significantly lower than those trained on UNSW-NB15 and BoT-IoT.
@@
It is observable that the KAN outperforms the MLP networks, which are tailored to have the same number of parameters as the KAN. In the end-to-end training paradigm, the networks are trained using only the fine-tuning set, without pre-training on unlabeled data. The contrastive training, by leveraging large volumes of unlabeled data, further improves the performance.
@@
It is worth noting that, for KANs, the width of the network has little impact on the performance. As can be observed, the networks with widths of 50 or 100 outperform those with widths of 150 and 200. This supports the earlier findings that the main factors influencing the performance of KAN are the grid size and spline order. However, for MLP networks, increasing the width often results in marginally better performance. This indicates that, for MLP networks, there is a more predictable relationship between network capacity and performance; in other words, MLP networks are easier to tune.
@@
Table 10: Comparison of our proposed methods with other approaches on the UNSW-NB15 dataset.
@@
Table 11: Comparison of our proposed methods with other approaches on the BoT-IoT dataset.
@@ 5.5 Comparative Analysis
Table 10, Table 11, and Table 12 present a comparative analysis of state-of-the-art methods. The performance of the proposed method is superior to other learning-based algorithms, some of which utilize more labeled data than our approach. This demonstrates that the proposed algorithm is well-suited for scenarios with label scarcity.
@@ 6 Conclusion
In this paper, we proposed a semi-supervised contrastive learning approach for real-time intrusion detection under limited labeled data conditions. This framework demonstrates strong performance on IoT and IIoT datasets, outperforming other state-of-the-art methods. The results of this study show that the performance of the KAN is superior to that of MLP networks when labeled data is scarce. Additionally, contrastive training improves the network’s performance by utilizing the available unlabeled data.
A key advantage of the proposed method is the interpretability provided by KAN, in the form of spline function visualization. This feature can be utilized to extract relevant rules from the network, which is beneficial in safety-critical applications, further increasing reliability, stability, and ensuring safe operation.
One of the findings of this paper is that the spline order and grid size are the most influential KAN hyper-parameters to tune, while the layer dimensions have minimal impact on improving performance. For our datasets, a higher grid size and lower spline order yield better results. Another key finding is that hyper-parameter tuning for MLP networks is much easier compared to their KAN counterparts, since KAN does not exhibit a straightforward relationship between performance and the number of parameters or the increase in spline order and grid size.
Although this study performed well on intrusion detection problems with limited labeled data, further studies must be conducted on the KAN architecture for the unsupervised case, where training requires no labeled data. This is left for future work.